### v1

@@ 5 Institute of Astrophysics, Foundation for Research and Technology - Hellas, Heraklion, GR7110, Greece
Evidence for a toroidal magnetic field in the core of 3C 84
The spatial scales of relativistic radio jets, probed by relativistic magneto-hydrodynamic jet launching simulations (RMHDs) and by most very-long-baseline interferometry (VLBI) observations differ by an order of magnitude. Bridging the gap between these RMHD simulations and VLBI observations requires selecting nearby active galactic nuclei (AGN), the parsec-scale region of which can be resolved. 3C 84 is a nearby bright AGN fulfilling the necessary requirements: it is launching a powerful, relativistic jet powered by a central supermassive black hole, while also being very bright. Using 22 GHz global VLBI measurements of 3C 84 we aim to study its sub-parsec region in both total intensity and linear polarisation, to explore the properties of this jet, with a linear resolution of parsec. We test different simulation setups by altering the bulk Lorentz factor of the jet, as well as the magnetic field configuration (toroidal, poloidal, helical). We confirm the persistence of a limb brightened structure, which reaches deep into the sub-parsec region. The corresponding electric vector position angles (EVPAs) follow the bulk jet flow inside but tend to be orthogonal to it near the edges. Our state-of-the-art RMHD simulations show that this geometry is consistent with a spine-sheath model, associated with a mildly relativistic flow and a toroidal magnetic field configuration.
@@
Making images of supermassive black holes’
citep[SMBHs;][]EHT19a, EHT22a, EHT24 shadows and their associated relativistic jets (Lu et al., 2023) have significantly improved our understanding of how these jets are launched.
Such a direct approach has only been possible for a select few radio sources; for all other sources, polarimetric observations are the key to study energy extraction from black holes.
SMBHs are thought to be capable of launching jets, both via its ergosphere (Blandford & Znajek, 1977, BZ model) and via its accretion disc (Blandford & Payne, 1982, BP model).
A stratified combination (Sol et al., 1989; Laing, 1996) of the two mechanisms can result in a spine-sheath type of flow (see also Hirotani et al., 2024).
This type of geometry has been revealed in a number of jets (see, for example, Attridge et al., 1999; Gabuzda et al., 2000; Pushkarev et al., 2005; O’Sullivan & Gabuzda, 2009; Mertens et al., 2016; Walker et al., 2018; Janssen et al., 2021; Lu et al., 2023) and attempts have been made to model it (Laing, 1996; Murphy et al., 2013; Laing & Bridle, 2014; Fuentes et al., 2018).
@@
The relativistic plasma and magnetic fields in these AGN jets generate incoherent synchrotron emission across a range of wavelengths, from radio to optical, UV, and possibly X-rays, characterised by both linear and circular polarisation. These polarisation characteristics (linear polarisation can reach values as high as 60-70% of the total intensity flux, as noted by Rybicki & Lightman 1979) provide insights into the jet’s physical properties, such as magnetic field intensity and configuration (Tsunetoe et al., 2020), structural layout, particle concentration, and plasma makeup, as detailed by Wardle et al. (1998).
@@
Images obtained in polarised emission are useful probes of jet physics because, in the non-relativistic case, the electric vector position angle (EVPA) is expected to align perpendicularly (in the optically thin regime) to the local magnetic field, aiding in understanding the magnetic field’s structure within the source (e.g., Traianou et al., 2024). In the case of the spine-sheath model, the EVPAs are oriented along the spine and perpendicular to the sheath. Relativistic magneto-hydrodynamic simulations (see e.g. Fuentes et al., 2018; Kramer & MacDonald, 2021) indicate that toroidal and helical magnetic field configurations are capable of producing such an EVPA configuration. In this case the envelope of the magnetic field lines in the sheath corresponds to the observed poloidal component, whereas the loops correspond to the toroidal component.
@@
Furthermore, using recent observations from the Imaging X-ray Polarimetry Explorer (IXPE, Weisskopf et al., 2022) Liodakis et al. (2022) have found that the electric vector polarisation angle of the millimetre-radio to X-ray emission is typically aligned with the jet axis. This was found to be true for the optically thin synchrotron emission in all sub-classes of blazars (e.g., Ehlert et al., 2023; Middei et al., 2023; Peirson et al., 2023). The IXPE observations have also discovered a rapid rotation of the polarisation angle in X-rays without an optical or radio counterpart (Di Gesu et al., 2023). The lack of a counterpart at lower frequencies points to non co-spatial emission regions, which is consistent with energy-stratified emission from a shock-in-jet scenario (Marscher & Gear, 1985).
@@
At a luminosity distance of 78.9 Mpc (, Strauss et al. 1992) in the Perseus galaxy cluster, 3C 84 (NGC 1275) is one of the nearest and brightest radio galaxies in the northern sky and thus a prime target to study jet formation. Its parsec-scale jet is characterised by limb brightening (Nagai et al., 2014; Giovannini et al., 2018), reaching deep into the sub-milliarcsecond (sub-mas) and contributing to its complex morphology (Hodgson et al., 2018; Paraschos et al., 2021; Punsly et al., 2021; Oh et al., 2022). Polarisation studies of the source have been carried out in the past (Homan & Wardle, 2004; Taylor et al., 2006; Plambeck et al., 2014; Nagai et al., 2017; Kim et al., 2019; Paraschos et al., 2024), providing insights into the interaction of the jet with the ambient medium, as well as the accretion state of the central SMBH.
@@
3C 84 has been monitored at centimetre wavelengths for decades (e.g. Marr et al., 1989; Vermeulen et al., 1994; Walker et al., 1994; Dhawan et al., 1998; Walker et al., 2000; Giovannini et al., 2018; Britzen et al., 2019; Paraschos et al., 2022; Savolainen et al., 2023) to study the parsec-scale jet structure111At greater distances from the central engine, on the decaparsec to kiloparsec-scale, even more structure is detected (Walker et al., 2000; Fabian et al., 2000).. There, three main areas of the approaching (Southern) jet can be distinguished (e.g. Nagai et al., 2014); they are usually labelled as ‘C1’ (‘core’), ‘C2’ (diffuse emission region Southwest of C1), and ‘C3’ (a prominent, moving hotspot South of C1, associated with the restarted jet activity; Nagai et al. 2010). Polarisation has been detected in both the compact sub-mas region and the extended, parsec-scale southern emission, although more prominently in the latter. Specifically, the parsec-scale region has been reported to exhibit circular polarisation (Homan & Wardle, 2004) and substantial Faraday rotation measures (Taylor et al., 2006). Newer measurements have shown that linear polarisation is also detected in C3 (Nagai et al., 2017) and in the ultimate vicinity of the central engine (Paraschos et al., 2024). Still, the polarisation signature in the core region of 3C 84 (C1) remains elusive, on distances of from the central engine.
@@
We thus utilised the excellent -coverage and sensitivity offered by global VLBI to examine the core region of 3C 84 in both total and polarised intensity. We then compared our observations with state-of-the-art RMHD simulations to draw conclusions about the magnetic field configuration in that core region.
@@
This letter is structured as follows: in Sect. 2 we briefly discuss our observations, data analysis, and results. In Sect. 3 we compare the resulting image to RMHD simulations and draw conclusions about the magnetic field configuration. Finally, in Sect. 4 we present our conclusions. Throughout the paper we assume a cold dark matter cosmology with , , and (Planck Collaboration et al., 2016).
@@ 2 Data, analysis, and results
3C 84 was observed in November 2021 with the European VLBI Network (EVN) in conjunction with further telescopes to form a global array, at both 22 and 43 GHz, as part of a multi-wavelength effort to study its parsec-scale jet via direct imaging and polarisation. In total twenty-two antennas successfully observed the source (they are listed in Appendix A.1). Calibration and imaging of the data was performed using the rPICARD pipeline (Janssen et al., 2019), the polsolve software (Martí-Vidal et al., 2021), and the difmap software (Shepherd et al., 1994). Details of the post-correlation, calibration, and imaging procedures are discussed in Appendices A.2 and A.3.
Here we present the 22 GHz image of 3C 84, as observed in Stokes and linear polarisation (see Fig. 1). Our observations constitute some of the highest resolution linearly polarised flux detections in AGN. Similar observations in sensitivity and -coverage were conducted as part the Global VLBI Alliance forum (GVA222http://gvlbi.evlbi.org/; Park et al., 2023) a year later and exhibit very similar structure in total intensity. That work also provides the most recent estimate of the bulk jet flow direction; they show that the jet propagates southeastwards in the compact core region, before bending southwards further downstream.
We find evidence of linearly polarised flux in the sub-mas region, tracing two limbs, the EVPAs of which are perpendicular to the bulk jet flow (regions ‘R’ and ‘R’ in Fig. 1). Specifically, region R is detected at a significance level exceeding in polarisation, whereas region R corresponds to a detection. This points to the prevalence of magnetic field lines running parallel to the direction of the southern jet propagation. Weak polarised flux is also detected between these two limbs, denoted as region ‘R’. Interestingly, the EVPAs seem transverse to the ones in the two limbs, indicating that the magnetic field there is perpendicular to the bulk jet flow and possibly indicative of a shock. A fourth region (‘R’) with polarised flux is also detected in the vicinity of where the jet curves southwards. We discuss the implications of our findings and compare them to simulations in the next section.
@@ 3.1 Spine-sheath geometry
Diverse mechanisms can give rise to a limb-brightened geometry; for example, a velocity gradient in the jet caused by shear acceleration at its boundary, jet stratification, geometrical projection effects, as well as the direct influence of the SMBH powering the jet (see, e.g. Blandford et al., 2019, for a relevant review). In the latter case, such a geometry, manifested in both Stokes (see Fig. 4) and linear polarisation, has implications on the conditions of the central engine (i.e. the nature of the accretion flow, the spin of the BH, and the magnetic state), as well as the magnetic field configuration downstream, in the sub-parsec region. To investigate the sub-milliarcsecond jet region, Takahashi et al. (2018) performed a simulation of a steady, axisymmetric force-free jet; they found that a limb-brightened, symmetric structure may be associated with an advection-dominated accretion flow (ADAF) and a magnetically arrested disc (MAD) around a rapidly spinning SMBH. Their findings were confirmed observationally, using the Event Horizon Telescope (EHT), to resolve the core region of 3C 84 at 230 GHz (see Paraschos et al., 2024).
@@
On the other hand, in the downstream region, helical and toroidal magnetic field lines have been associated with an increase in linear polarisation towards the jet edges (Pushkarev et al., 2005; Lyutikov et al., 2005). Specifically, it is expected that toroidal magnetic field geometries give rise to symmetric Stokes and linear polarisation profiles, whereas helical ones cause asymmetric ones (Gabuzda, 2020; Kramer & MacDonald, 2021). In addition to the intrinsic magnetic field carried by the jet, shear from velocity gradients in the flow might also contribute to the observed geometry (e.g. Laing, 1980). The EVPAs associated with these profiles are expected to be bimodal, that is, perpendicular to the bulk jet flow on the bright limbs and parallel to it on the inner ridge (O’Sullivan & Gabuzda, 2009, see also Kramer & MacDonald 2021 for RMHD simulations revealing the EVPA orientation for different magnetic field geometries).
@@
Here we adapted these RMHD simulations to the parameters of 3C 84. A more detailed description of the simulations setup is described in Appendix B.1. Specifically, we assume a luminosity distance of and investigate the effects of different bulk Lorentz factors and magnetic field configurations (toroidal, poloidal, helical; detailed descriptions and illustrations of the magnetic field vectors can be found in Kramer & MacDonald, 2021). For the fiducial simulation presented in Fig. 2, we chose a bulk Lorentz factor of (the slow pattern speeds in the sub-mas region of 3C 84, e.g. Krichbaum et al., 1992; Dhawan et al., 1998; Hodgson et al., 2021; Weaver et al., 2022; Paraschos et al., 2022, suggest that ) and a toroidal magnetic field configuration. The figure shows linearly polarised intensity () over-plotted with both total intensity contours and EVPAs (). The integrated level of fractional linear polarisation is listed to the lower left, that is (barred quantities correspond to flux weighted averages). In Appendix B.2 we present alternative simulation results, based on these other parameters. The uncertainty of the SMBH’s mass in 3C 84 (mass ranges between have been reported in previous studies, e.g. Park & Trippe 2017; Giovannini et al. 2018) does not affect our simulations, given that general relativistic effects should be negligible at these distances from the central engine.
@@
We find that the geometry of the EVPAs on the simulated double rail structure corresponds to a toroidal magnetic field geometry; this is also seen in our observations in Fig. 1. The dominance of the toroidal magnetic field component beyond the recollimation region (Laing & Bridle, 2014) is also a natural consequence of jet expansion, as the toroidal component’s strength decreases slower than that of the poloidal component (Burch, 1979; Fuentes et al., 2018). This magnetic field geometry, as indicated by our observations, could imply a connection to magnetic jet launching, as is the case for the BZ model. Our finding is in line with mounting evidence connecting the jet launching in 3C 84 to the BZ model (e.g. Takahashi et al., 2018; Oh et al., 2022; Paraschos et al., 2023, 2024).
@@
In light of the recent IXPE observations of blazars, our results also suggest that the spine dominates the radio to X-ray synchrotron emission, where electron populations are accelerated in shocks, hence the alignment of the polarisation vectors with the jet. The emission then becomes energy-stratified due to particle cooling, as the particles propagate downstream from the shock front. This would also suggest the X-ray polarisation angle rotation was possibly the result a shock moving through the helical magnetic field (see proposed model discussion in Kim et al., 2024).

### v2

@@ 5 Institute of Astrophysics, Foundation for Research and Technology - Hellas, Heraklion, GR7110, Greece
Evidence of a toroidal magnetic field in the core of 3C 84
The spatial scales of relativistic radio jets, probed by relativistic magneto-hydrodynamic (RMHD) jet launching simulations and by most very long baseline interferometry (VLBI) observations differ by an order of magnitude. Bridging the gap between these RMHD simulations and VLBI observations requires selecting nearby active galactic nuclei (AGN), the parsec-scale region of which can be resolved. The radio source 3C 84 is a nearby bright AGN fulfilling the necessary requirements: it is launching a powerful, relativistic jet powered by a central supermassive black hole, while also being very bright. Using 22 GHz globe-spanning VLBI measurements of 3C 84 we studied its sub-parsec region in both total intensity and linear polarisation to explore the properties of this jet, with a linear resolution of parsec. We tested different simulation set-ups by altering the bulk Lorentz factor of the jet, as well as the magnetic field configuration (toroidal, poloidal, helical). We confirm the persistence of a limb brightened structure, which reaches deep into the sub-parsec region. The corresponding electric vector position angles (EVPAs) follow the bulk jet flow inside but tend to be orthogonal to it near the edges. Our state-of-the-art RMHD simulations show that this geometry is consistent with a spine-sheath model, associated with a mildly relativistic flow and a toroidal magnetic field configuration.
@@
Making images of supermassive black hole (SMBH; EHTC et al., 2019a, 2022a, 2024) shadows and their associated relativistic jets (Lu et al., 2023) have significantly improved our understanding of how these jets are launched. Such a direct approach has only been possible for a select few radio sources; for all other sources, polarimetric observations are key for studying energy extraction from black holes. A SMBH is thought to be capable of launching jets, both via its ergosphere (Blandford & Znajek, 1977, BZ model) and via its accretion disc (Blandford & Payne, 1982, BP model). A stratified combination (Sol et al., 1989; Laing, 1996) of the two mechanisms can result in a spine-sheath type of flow (see also Hirotani et al., 2024). This type of geometry has been revealed in a number of jets (see e.g. Attridge et al., 1999; Gabuzda et al., 2000; Pushkarev et al., 2005; O’Sullivan & Gabuzda, 2009; Mertens et al., 2016; Walker et al., 2018; Janssen et al., 2021; Lu et al., 2023) and attempts have been made to model it (Laing, 1996; Murphy et al., 2013; Laing & Bridle, 2014; Fuentes et al., 2018).
@@
The relativistic plasma and magnetic fields in these AGN jets generate incoherent synchrotron emission across a range of wavelengths, from radio to optical, UV, and possibly X-rays, characterised by both linear and circular polarisation. Linear polarisation can reach values as high as 60-70% of the total intensity flux, as noted by Rybicki & Lightman 1979. These polarisation characteristics provide insights into the jet’s physical properties, such as magnetic field intensity and configuration (Tsunetoe et al., 2020), structural layout, particle concentration, and plasma makeup, as detailed by Wardle et al. (1998).
@@
Images obtained in polarised emission are useful probes of jet physics because, in the non-relativistic case, the electric vector position angle (EVPA) is expected to align perpendicularly (in the optically thin regime) to the local magnetic field, aiding in understanding the magnetic field’s structure within the source (e.g. Traianou et al., 2024). In the case of the spine-sheath model, the EVPAs are oriented along the spine and perpendicular to the sheath. Relativistic magneto-hydrodynamic simulations (see e.g. Fuentes et al., 2018; Kramer & MacDonald, 2021) indicate that toroidal and helical magnetic field configurations are capable of producing such an EVPA configuration. In this case the envelope of the magnetic field lines in the sheath corresponds to the observed poloidal component, whereas the loops correspond to the toroidal component.
@@
Furthermore, using recent observations from the Imaging X-ray Polarimetry Explorer (IXPE, Weisskopf et al., 2022) Liodakis et al. (2022) have found that the electric vector polarisation angle of the millimetre-radio to X-ray emission is typically aligned with the jet axis. This was found to be true for the optically thin synchrotron emission in all sub-classes of blazars (e.g. Ehlert et al., 2023; Middei et al., 2023; Peirson et al., 2023). The IXPE observations have also discovered a rapid rotation of the polarisation angle in X-rays without an optical or radio counterpart (Di Gesu et al., 2023). The lack of a counterpart at lower frequencies points to non-cospatial emission regions, which is consistent with energy-stratified emission from a shock-in-jet scenario (Marscher & Gear, 1985).
@@
At a luminosity distance of 78.9 Mpc (, Strauss et al. 1992) in the Perseus galaxy cluster, 3C 84 (NGC 1275) is one of the nearest and brightest radio galaxies in the northern sky, and thus is a prime target to study jet formation. Its parsec-scale jet is characterised by limb brightening (Nagai et al., 2014; Giovannini et al., 2018), reaching deep into the sub-milliarcsecond (sub-mas) and contributing to its complex morphology (Hodgson et al., 2018; Paraschos et al., 2021; Punsly et al., 2021; Oh et al., 2022). Polarisation studies of the source have been carried out in the past (Homan & Wardle, 2004; Taylor et al., 2006; Plambeck et al., 2014; Nagai et al., 2017; Kim et al., 2019; Paraschos et al., 2024), providing insights into the interaction of the jet with the ambient medium, as well as the accretion state of the central SMBH.
@@
The radio source 3C 84 has been monitored at centimetre wavelengths for decades (e.g. Marr et al., 1989; Vermeulen et al., 1994; Walker et al., 1994; Dhawan et al., 1998; Walker et al., 2000; Giovannini et al., 2018; Britzen et al., 2019; Paraschos et al., 2022; Savolainen et al., 2023) to study the parsec-scale jet structure111At greater distances from the central engine, on the decaparsec to kiloparsec scales, even more structure is detected (Walker et al., 2000; Fabian et al., 2000).. Three main areas of the approaching (southern) jet can be distinguished (e.g. Nagai et al., 2014); they are usually labelled as C1 (core), C2 (diffuse emission region south-west of C1), and C3 (a prominent, moving hotspot south of C1, associated with the restarted jet activity; Nagai et al. 2010). Polarisation has been detected in the compact sub-mas region and in the extended parsec-scale southern emission, although more prominently in the latter. Specifically, the parsec-scale region has been reported to exhibit circular polarisation (Homan & Wardle, 2004) and substantial Faraday rotation measures (Taylor et al., 2006). Newer measurements have shown that linear polarisation is also detected in C3 (Nagai et al., 2017) and in the ultimate vicinity of the central engine (Paraschos et al., 2024). Still, the polarisation signature in the core region of 3C 84 (C1) remains elusive, at distances of from the central engine.
@@
We thus utilised the excellent -coverage and sensitivity offered by globe-spanning very long baseline interferometry (VLBI) to examine the core region of 3C 84 in both total and polarised intensity. We then compared our observations with state-of-the-art RMHD simulations to draw conclusions about the magnetic field configuration in that core region.
@@
This letter is structured as follows. In Sect. 2 we briefly discuss our observations, data analysis, and results. In Sect. 3 we compare the resulting image to RMHD simulations and draw conclusions about the magnetic field configuration. Finally, in Sect. 4 we present our conclusions. Throughout the paper we assume a cold dark matter cosmology with , , and (Planck Collaboration et al., 2016).
@@ 2 Data, analysis, and results
The radio source 3C 84 was observed in November 2021 with the European VLBI Network (EVN) in conjunction with other telescopes, to form a globe-spanning array, at both 22 and 43 GHz, as part of a multi-wavelength effort to study its parsec-scale jet via direct imaging and polarisation. In total, 22 antennas successfully observed the source (see Appendix A.1). Calibration and imaging of the data were performed using the rPICARD pipeline (Janssen et al., 2019), the polsolve software (Martí-Vidal et al., 2021), and the difmap software (Shepherd et al., 1994). Details of the post-correlation, calibration, and imaging procedures are discussed in Appendices A.2 and A.3.
Here we present the 22 GHz image of 3C 84, as observed in Stokes and linear polarisation (see Fig. 1). Our observations constitute some of the highest resolution linearly polarised flux detections in AGN. Similar observations in sensitivity and -coverage were conducted as part the Global VLBI Alliance forum (GVA;222http://gvlbi.evlbi.org/ Park et al., 2023) a year later, and exhibit very similar structure in total intensity. That work also provides the most recent estimate of the bulk jet flow direction; the authors show that the jet propagates south-eastwards in the compact core region, before bending southwards further downstream.
We find evidence of linearly polarised flux in the sub-mas region, tracing two limbs, the EVPAs of which are perpendicular to the bulk jet flow (regions R and R in Fig. 1). Specifically, region R is detected at a significance level exceeding in polarisation, whereas region R corresponds to a detection. This points to the prevalence of magnetic field lines running parallel to the direction of the southern jet propagation. Weak polarised flux is also detected between these two limbs, denoted as region R. Interestingly, the EVPAs seem transverse to those in the two limbs, indicating that the magnetic field there is perpendicular to the bulk jet flow and possibly indicative of a shock. A fourth region (R) with polarised flux is also detected in the vicinity of where the jet curves southwards. We discuss the implications of our findings and compare them to simulations in the next section.
@@ 3.1 Spine-sheath geometry
Diverse mechanisms can give rise to a limb-brightened geometry, for example a velocity gradient in the jet caused by shear acceleration at its boundary, jet stratification, geometrical projection effects, as well as the direct influence of the SMBH powering the jet (see e.g. Blandford et al., 2019, for a relevant review). In the latter case, this geometry, manifested in both Stokes (see Fig. 4) and linear polarisation, has implications for the conditions of the central engine (i.e. the nature of the accretion flow, the spin of the BH, and the magnetic state) and for the magnetic field configuration downstream, in the sub-parsec region. To investigate the sub-milliarcsecond jet region, Takahashi et al. (2018) performed a simulation of a steady, axisymmetric force-free jet. They found that a limb-brightened symmetric structure may be associated with an advection-dominated accretion flow (ADAF) and a magnetically arrested disc (MAD) around a rapidly spinning SMBH. Their findings were confirmed observationally, using the Event Horizon Telescope (EHT), to resolve the core region of 3C 84 at 230 GHz (see Paraschos et al., 2024).
@@
On the other hand, in the downstream region, helical and toroidal magnetic field lines have been associated with an increase in linear polarisation towards the jet edges (Pushkarev et al., 2005; Lyutikov et al., 2005). Specifically, it is expected that toroidal magnetic field geometries give rise to symmetric Stokes and linear polarisation profiles, whereas helical geometries cause asymmetric profiles (Gabuzda, 2020; Kramer & MacDonald, 2021). In addition to the intrinsic magnetic field carried by the jet, shear from velocity gradients in the flow might also contribute to the observed geometry (e.g. Laing, 1980). The EVPAs associated with these profiles are expected to be bimodal, that is, perpendicular to the bulk jet flow on the bright limbs and parallel to it on the inner ridge (O’Sullivan & Gabuzda, 2009, see also Kramer & MacDonald 2021 for RMHD simulations revealing the EVPA orientation for different magnetic field geometries).
@@
Here we adapted these RMHD simulations to the parameters of 3C 84. A more detailed description of the simulations set-up is described in Appendix B.1. Specifically, we assume a luminosity distance of and investigate the effects of different bulk Lorentz factors and magnetic field configurations (toroidal, poloidal, helical; detailed descriptions and illustrations of the magnetic field vectors can be found in Kramer & MacDonald, 2021). For the fiducial simulation presented in Fig. 2, we chose a bulk Lorentz factor of (the slow pattern speeds in the sub-mas region of 3C 84, e.g. Krichbaum et al., 1992; Dhawan et al., 1998; Hodgson et al., 2021; Weaver et al., 2022; Paraschos et al., 2022, suggest that ) and a toroidal magnetic field configuration. The figure shows linearly polarised intensity () overplotted with both total intensity contours and EVPAs (). The integrated level of fractional linear polarisation is shown in the lower left, that is (barred quantities correspond to flux weighted averages). In Appendix B.2 we present alternative simulation results, based on these other parameters. The uncertainty of the SMBH mass in 3C 84 (mass ranges between have been reported in previous studies, e.g. Park & Trippe 2017; Giovannini et al. 2018) does not affect our simulations, given that general relativistic effects should be negligible at these distances from the central engine.
@@
We find that the geometry of the EVPAs on the simulated double rail structure corresponds to a toroidal magnetic field geometry; this is also seen in our observations in Fig. 1. The dominance of the toroidal magnetic field component beyond the recollimation region (Laing & Bridle, 2014) is also a natural consequence of jet expansion, as the toroidal component’s strength decreases more slowly than that of the poloidal component (Burch, 1979; Fuentes et al., 2018). This magnetic field geometry, as indicated by our observations, could imply a connection to magnetic jet launching, as is the case for the BZ model. Our finding is in line with mounting evidence connecting the jet launching in 3C 84 to the BZ model (e.g. Takahashi et al., 2018; Oh et al., 2022; Paraschos et al., 2023, 2024).
@@
In light of the recent IXPE observations of blazars, our results also suggest that the spine dominates the radio to X-ray synchrotron emission, where electron populations are accelerated in shocks, hence the alignment of the polarisation vectors with the jet. The emission then becomes energy-stratified due to particle cooling, as the particles propagate downstream from the shock front. This would also suggest the X-ray polarisation angle rotation was possibly the result of a shock moving through the helical magnetic field (see the discussion on the proposed model in Kim et al., 2024).